Solve the system of equations
x + 3y = 32
-3x + 2y = 3

Answer by mananth(16949)

1 x + 3 y = 32 .............1
-3 x + 2 y = 3 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 3
-2 x -6 y = -64
-9 x 6 y = 9
Add the two equations
-11 x = -55
/ -11
x = 5
plug value of x in (1)
1 x + 3 y = 32
5 + 3 y = 32
3 y = 32 -5
3 y = 27
y = 9